one side of a rectangular stage is 2 meters longer than the other. if the diagonal is 10 meters, then what are the length of the sides?
Answer by user_dude2008(1862)

x^2+(x+2)^2=10^2


x^2+x^2+4x+4=100

2x^2+4x-96=0

x^2+2x-48=0

(x-6)(x+8)=0

x-6=0 or x+8=0

x=6 or x=8

Answer: side lengths are 6 and 8